OpenStudy (anonymous):

Simplify: (6/x-6)+(3x/x+5)-1/(x+6)(x+5)

OpenStudy (anonymous):

Please help

OpenStudy (tkhunny):

The first term is \(\dfrac{6}{x} - 6\) but you probably meant \(\dfrac{6}{x-6}\) The second term is \(\dfrac{3x}{x} + 6\) but you probably meant \(\dfrac{3x}{x+5}\) The third term is \(\dfrac{1}{x+6}\cdot (x+5)\) but you probably meant \(\dfrac{1}{(x+6)(x+5)}\). Come to think of it, you probably meant \(\dfrac{1}{(x-6)(x+5)}\) So, please type in teh problem that you mean to present. Then, use your best fraction-adding skills to give it a go!

OpenStudy (anonymous):

Do you know how I could solve this? @tkhunny

OpenStudy (tkhunny):

You did not understand my last post. I do not yet know what the problem is. Also, I suspect it is just adding fractions. Do you remember how to do it? \(\dfrac{2}{3} + \dfrac{4}{5} = \dfrac{2}{3}\dfrac{5}{5} + \dfrac{4}{5}\dfrac{3}{3} = \dfrac{10}{15} + \dfrac{12}{15} = \dfrac{22}{15}\) It is just the same with binomials.